OpenStudy (anonymous):

Determine the equation of the line passing through (2, -3) and parallel to the line passing through (4,1) and (-2,2).

OpenStudy (mathmate):

Hint: 1. find the slope of the line passing through (4,1) and (-2,2) using slope, a=(y2-y1)/(x2-x1) 2. find the line passing through (x0,y0) using y=a(x-x0)+y0

OpenStudy (anonymous):

two lines are parallel when their slopes are equal

OpenStudy (anonymous):

and the slope of a line is (y2-y1)/(x2-x1)

OpenStudy (anonymous):

slope of second line= 2-1/-2-4 =1/-6 =-1/6

OpenStudy (anonymous):

so by using point slope form,(y--3)= -1/6(x-2)

OpenStudy (anonymous):

Ohhh. I get it. Well thank you @mathmate and @krishnadas for your help! :)